The perimeter of a rectangle is 58 centimeters . The length of the rectangle is four less than twice the width of the rectangle. Find the length, the width, and the area of this rectangle.
Answer by TimothyLamb(4379)

---
2x + 2y = 58
x = 2y - 4
---
put the system of linear equations into standard form:
---
2x + 2y = 58
x - 2y = -4
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x length = 18 cm
y width = 11 cm
---
system matrix
2 2
1 -2
---
inverse of system matrix
0.33333333 0.33333333
0.16666667 -0.33333333
---
determinant of system matrix = -6
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php